### v1

# Subgraphs with a positive minimum semidegree in digraphs with large outdegree

Abstract
We prove that every $n$-vertex directed graph $G$ with the minimum outdegree $\delta^{+}(G)=d$ contains a subgraph $H$ satisfying
$\min\left\{\delta^{+}(H),\delta^{-}(H)\right\}\geq\frac{d(d+1)}{2n}\,.$
We also show that if $d=o(n)$ then this bound is asymptotically best possible.
1 Introduction
A well known result of Erdős [5] states that every undirected graph with $n$ vertices and $m$ edges contains a subgraph with minimum degree at least $\frac{m}{n}$. In order to establish a meaningful analogue of this statement in the setting of directed graphs, one needs to impose some additional assumptions; indeed every subgraph of a transitive tournament contains a vertex with zero indegree (and similarly a vertex with zero outdegree). Huang et al. [6] proved that if a directed graph with $n$ vertices and $m$ arcs is Eulerian, i.e., if every vertex has its indegree equal to its outdegree, then one can always find an Eulerian subgraph with the minimum indegree at least $\frac{m^{2}}{24n^{3}}$.
In this paper, we focus on directed graphs where every vertex has a large outdegree. It is easy to observe that if every vertex in a finite digraph has non-zero outdegree, then it has to contain a directed cycle. An immediate corollary of this observation is that such a digraph contains a subgraph where both the minimum indegree and the minimum outdegree are non-zero. Our aim is to study a quantitative strengthening of the aforementioned corollary.
1.1 Notation
Given a finite vertex-set $V$, a directed graph, or digraph for short, is a pair $(V,A)$, where $A\subseteq V\times V$. The elements of $A$ are called arcs, and for a given arc $e=(u,w)$, where $u,w\in V$, we say that $e$ is oriented from $u$ to $w$. We note that all the digraphs in this paper are loopless, i.e., $A\subseteq V\times V\setminus\{(v,v):v\in V\}$, however, pairs of different vertices are allowed to be connected by two arcs with the opposite directions. For a systematic study of digraphs, we refer the reader to the monograph of Bang-Jensen and Gutin [2].
Given a digraph $G=(V,A)$ and a vertex $v\in V$, we define the outneighborhood of $v$ and inneighborhood of $v$ to be
$N_{G}^{+}(v):=\{w\in V:(v,w)\in A\}\quad\mbox{and}\quad N_{G}^{-}(v):=\{u\in V:(u,v)\in A\}\,,$
respectively. We set $\operatorname{deg}^{+}_{G}(v):=|N^{+}_{G}(v)|$ and $\operatorname{deg}^{-}_{G}(v):=|N_{G}^{-}(v)|$, and call the respective quantities the outdegree of $v$ and the indegree of $v$. We drop the subscript in cases when the digraph $G$ is clear from the context.
For a digraph $G$ on the vertex-set $V$, we define the minimum outdegree of $G$, which we denote by $\delta^{+}(G)$, to be the smallest outedgree of a vertex in $G$. In other words, $\delta^{+}(G):=\min_{v\in V}\operatorname{deg}^{+}_{G}(v)$. Analogously, we define the minimum indegree of $G$ as $\delta^{-}(G):=\min_{v\in V}\operatorname{deg}^{-}_{G}(v)$. Finally, the minimum semidegree of $G$, which we denote by $\delta^{\pm}(G)$, is the minimum of $\delta^{+}(G)$ and $\delta^{-}(G)$.
1.2 Our results
Using the notation from Section 1.1, the observation from the beginning can be reformulated as follows: every digraph $G$ with non-zero $\delta^{+}(G)$ contains a subgraph $H$ with $\delta^{\pm}(H)\geq 1$. It is natural to ask whether a stronger lower bound on $\delta^{+}(G)$ guarantees a stronger bound on $\delta^{\pm}(H)$. The following example shows that even an assumption $\delta^{+}(G)=\sqrt{2n}-2$ is not sufficient to guarantee a subgraph $H$ in $G$ with $\delta^{\pm}(H)\geq 2$:
  1. 1.
    Partition the set $[n]$ into two sets $A:=\left[\sqrt{2n}\right]$ and $B:=[n]\setminus A$.
  2. 2.
    Orient all the pairs $i<j$ in $A$ from $j$ to $i$.
  3. 3.
    For every $i\in A$, select $\sqrt{2n}-i-1$ outneighbors of $i$ in $B$ such a way that every $v\in B$ has at most one inneighbor in $A$; note that this can be done since $\sum_{i=1}^{\sqrt{2n}}\sqrt{2n}-i-1=n-\sqrt{4.5n}<|B|$.
  4. 4.
    Orient all the remaining pairs between $A$ and $B$ from $B$ to $A$.
A moment of thought reveals that every vertex has at least $\sqrt{2n}-2$ outneighbors, yet every subgraph $H$ satisfies $\delta^{\pm}(H)\leq 1$; see Section 3 for details. The main result of this paper is that a lower bound $\delta^{+}(G)\geq\sqrt{2n}$ already yields a subgraph $H$ with $\delta^{\pm}(H)>1$.
Theorem 1.
Every loopless digraph $G$ with $n$ vertices and $\delta^{+}(G)\geq d$ has a subgraph $H$ with $\delta^{\pm}(H)\geq\frac{d(d+1)}{2n}$.
The rest of the paper is organized as follows. We prove Theorem 1 in Section 2. In Section 3, we generalize the construction from the previous paragraph and show that Theorem 1 is for $d=o(n)$, up to lower order terms, best possible. We conclude by Section 4, where we discuss the situation for the dense case $d=\alpha n$, and state some related open problems.
2 Proof of Theorem 1
Fix an $n$-vertex digraph $G$. Without loss of generality, we will assume that every vertex of $G$ has the outdegree equal to $d$, as otherwise we pass to an appropriate spanning subgraph of $G$. Now consider the following greedy procedure of one-by-one removing vertices from $G$:
  1. 1.
    Let $G_{n}:=G$.
  2. 2.
    For $i=n,n-1,\ldots,2,1$, do:
    1. (a)
      Let $v_{i}$ be a vertex in $G_{i}$ with $\min\{\deg_{G_{i}}^{+}(v_{i}),\deg_{G_{i}}^{-}(v_{i})\}=\delta^{\pm}(G_{i})$.
    2. (b)
      Set $G_{i-1}:=G_{i}-v_{i}$.
We note that in case of a tie between the minimum indegree and the minimum outdegree during the $i$-th step, i.e., when $\delta^{+}(G_{i})=\delta^{-}(G_{i})$, we choose as $v_{i}$ a vertex that has the smallest indegree.
For the rest of the proof, we will be considering the vertices of $G$ being ordered as $v_{1}<v_{2}<\ldots<v_{n}$. In other words, it is the reversed order with respect to the removal during the greedy procedure. For brevity, given a vertex $v\in V(G)$, we will be writing $N^{+}_{L}(v)$ and $N^{+}_{R}(v)$ to denote the set of the outneighbors of $v$ in $G$ that are in the order before $v$ and after $v$, respectively. Analogously, we define $N^{-}_{L}$ and $N^{-}_{R}$. In other words,
$N^{+}_{L}(v):=$$\left\{w\in N_{G}^{+}(v):w<v\right\}\,,$ $N^{+}_{R}(v):=$$\left\{w\in N_{G}^{+}(v):w>v\right\}\,,$
$N^{-}_{L}(v):=$$\left\{w\in N_{G}^{-}(v):w<v\right\}\,,$ $N^{-}_{R}(v):=$$\left\{w\in N_{G}^{-}(v):w>v\right\}\,.$
Let $V^{-}:=\left\{v_{i}:\operatorname{deg}^{-}_{G_{i}}(v_{i})=\delta^{\pm}(G_{i})\right\}$, and $V^{+}:=V(G)\setminus V^{-}$. In words, the set $V^{-}$ consists of all the vertices that the greedy procedure removed because of their small indegree, and the vertices in $V^{+}$ are those that were removed because of their small outdegree. Set $c:=\max_{i\in[n]}\delta^{\pm}(G_{i})$. Clearly, it suffices to show that $c\geq\frac{d(d+1)}{2n}$. We will do so by distinguishing two cases based on the size of $V^{+}$.
We first focus on the case $|V^{+}|\geq d$. Let $X$ be the last $d$ vertices from $V^{+}$, i.e., the first $d$ vertices that the greedy procedure removed because of having the smallest outdegree in the corresponding subgraph. Since every vertex of the digraph $G$ has outregee equal to $d$, it holds that $d^{2}=\sum_{x\in X}\operatorname{deg}_{G}^{+}(x)$. On the other hand, every vertex $x\in X$ satisfies
$\operatorname{deg}^{+}_{G}(x)=|N^{+}_{L}(x)|+|N^{+}_{R}(x)\cap X|+|N^{+}_{R}(x)\setminus X|\,.$
By the definition of $X\subseteq V^{+}$, every $x\in X$ also satisfies $|N^{+}_{L}(x)|\leq c$. Furthermore, we claim that
$\sum_{x\in X}|N^{+}_{R}(x)\cap X|\leq\binom{d}{2}\,.$ (1)
Indeed, every pair of the vertices from $X$ can contribute to this sum by at most one. Finally, the fact that $X$ consists of the rightmost vertices of $V^{+}$ implies that $\bigcup\limits_{x\in X}\left(N_{R}^{+}(x)\setminus X\right)\subseteq V^{-}$, which together with a standard double-counting argument yields that
$\sum_{x\in X}|N^{+}_{R}(x)\setminus X|=\sum_{x\in X}|N^{+}_{R}(x)\cap V^{-}|=\sum_{w\in V^{-}}|N^{-}_{L}(w)\cap X|\leq(n-d)c\,.$
Therefore, we conclude that
$d^{2}=\sum_{x\in X}|N^{+}_{L}(x)|+|N^{+}_{R}(x)\cap X|+|N^{+}_{R}(x)\setminus X|\leq d\cdot c+\binom{d}{2}+(n-d)c\,.$
Rearranging the last inequality readily yields that $c\geq\frac{d(d+1)}{2n}$.
It remains to consider the case when $|V^{+}|<d$. This time, let $Y^{-}$ be the first $\left(d-|V^{+}|\right)$ vertices from $V^{-}$, and let $Y:=Y^{-}\cup V^{+}$. Firstly, observe that all the vertices to the left of a vertex $y\in Y^{-}$ must lie inside the set $Y$. In particular, $N_{L}^{-}(y)\subseteq Y$ and $N_{L}^{+}(y)\subseteq Y$ for any $y\in Y^{-}$. Analogously as in (1), it holds that
$\sum_{y\in Y^{-}}|N^{+}_{L}(y)|+\sum_{y\in Y}|N^{+}_{R}(y)\cap V^{+}|\leq\binom{d}{2}\,,$
since the left-hand side of the inequality counts every pair of the vertices from $Y$ at most once. We also have that $|N^{+}_{L}(y)|\leq c$ for every $y\in V^{+}$, and additionally, it holds that $|N^{-}_{L}(y)|\leq c$ for every $y\in Y^{-}$. Thus,
$\sum_{y\in V^{+}}|N^{+}_{L}(y)|+\sum_{y\in Y}|N^{+}_{R}(y)\cap Y^{-}|\leq|V^{+}|\cdot c+\sum_{w\in Y^{-}}|N^{-}_{L}(w)|\leq d\cdot c\,.$
Finally, a double-counting argument yields that
$\sum_{y\in Y}|N^{+}_{R}(y)\setminus Y|=\sum_{w\in V\setminus Y}|N^{-}_{L}(w)\cap Y|\leq(n-d)c\,.$
Since
$|N_{R}^{+}(y)|=\left|N_{R}^{+}(y)\cap V^{+}\right|+\left|N_{R}^{+}(y)\cap Y^{-}\right|+\left|N_{R}^{+}(y)\setminus Y\right|\mbox{ for all }y\in Y,$
we have that
$d^{2}=\sum_{y\in Y}|N^{+}_{R}(y)|+\sum_{y\in Y^{-}}|N_{L}^{+}(y)|+\sum_{y\in V^{+}}|N_{L}^{+}(y)|\leq\binom{d}{2}+d\cdot c+(n-d)c.$
As in the previous case, rearranging $d^{2}\leq n\cdot c+\binom{d}{2}$ yields that $c\geq\frac{d(d+1)}{2n}$.
3 Tournaments with all subgraphs having small $\delta^{\pm}$
In this section, we prove that Theorem 1 is essentially best possible, even if we would consider only tournaments, i.e., orientations of the edge-set of complete undirected graphs.
Proposition 2.
Fix two positive integers $k$ and $\ell$, and let $d:=2k\ell-1$ and $n_{0}:=(k+1)d+\ell$. For every integer $n\geq n_{0}$ there exists a tournament $T$ with $n$ vertices and $\delta^{+}(T)=d$ such that
$\max_{H\subseteq T}\delta^{\pm}(H)\leq\ell\leq\left(1+\frac{k+1}{k^{2}}\right)\cdot\frac{d(d+1)}{2n}.$
Figure 1: The tournament constructed in Proposition 2.
Proof.
Fix the parameters $k$ and $\ell$. Firstly, let us consider the case $n=n_{0}$, and let $V$ be any vertex-set of cardinality $n$. We partition $V$ into three parts $A$, $B$ and $C$ of respective sizes $\ell$, $k\cdot d$ and $d$. Let us now describe the orientations of pairs of vertices from $V$. Firstly,
  1. 1.
    each of the sets $A$, $B$, $C$ induces a transitive orientation,
  2. 2.
    all the pairs between $A$ and $B$ are oriented from $B$ to $A$, and
  3. 3.
    all the pairs between $A$ and $C$ are oriented from $A$ to $C$.
Clearly, $\deg^{+}_{T}(v)\geq|C|=d$ for every vertex $v\in A$. Now, in order to have $\operatorname{deg}_{T}^{+}(v)=d$ for every $v\in C$, we need to orient a total of $\binom{d+1}{2}$ pairs from $C$ to $B$, where the $i$-th vertex of $C$ with respect to the transitive order needs to have exactly $i$ outneighbors in $B$. Our aim is to do this in such a way that every vertex of $B$ will have exactly $\frac{d+1}{2k}=\ell$ incoming arcs from $C$. We claim that this is always possible. Indeed, order $B$ arbitrarily and let $S$ be a sequence of length $\ell\cdot|B|$ formed by $\ell$ copies of the ordered $B$ placed after each other. Since $\ell|B|=\binom{d+1}{2}$ and $|B|\geq d$, we can go over the vertices in $C$ one by one and orient arcs from the $i$-th vertex to the next $i$ members of $S$. This establishes the desired $\operatorname{deg}_{T}^{+}(v)=d$ for every $v\in C$ as well as $|N^{-}(w)\cap C|=\ell$ for every $w\in B$. We orient all the remaining pairs between $B$ and $C$ from $B$ to $C$. In particular, every vertex in $B$ has now exactly $d-\ell+|A|=d$ outneighbors in $C\cup A$.
In the previous paragraph we have shown that $\delta^{+}(T)=d$, thus it remains to find an upper bound on $\delta^{\pm}(H)$ for any given $H\subseteq T$. Without loss of generality, we will assume that $H$ is induced. If $V(H)$ is disjoint from $B$ then $\delta^{+}(H)=\delta^{-}(H)=0$, because $A\cup C$ induces a transitive tournament. On the other hand, in case $V(H)\cap B\neq\emptyset$, let $v$ be the vertex with zero indegree in the transitive tournament induced by $V(H)\cap B$. Since $\deg^{-}_{H}(v)\leq\deg^{-}_{T}(v)=\ell$, we conclude that
$\delta^{\pm}(H)\leq\ell=\frac{d+1}{2k}=\left(1+\frac{d+\ell}{dk}\right)\cdot\frac{d(d+1)}{2n}\leq\left(1+\frac{k+1}{k^{2}}\right)\cdot\frac{d(d+1)}{2n}\,.$
It remains to consider the case $n>n_{0}$. Given the parameters $k$ and $\ell$, let $T_{0}$ be the previously constructed tournament on $n_{0}$ vertices. In order to construct the desired tournament $T$, add to $T_{0}$, one by one, a total of $n-n_{0}$ vertices in such a way that every newly added vertex has the indegree zero. The obtained tournament $T$ satisfies $\delta^{+}(T)=\delta^{+}(T_{0})=d$, and any $H\subseteq T$ with $\delta^{-}(H)>0$ satisfies that $H\subseteq T_{0}$. In particular, it holds that $\delta^{\pm}(H)\leq\ell$ for every $H\subseteq T$. ∎
4 Dense digraphs
If $d=o(n)$, then Theorem 1 and Proposition 2 determine, up to a multiplicative factor $(1+o(1))$, the best possible bound on $\max_{H\subseteq G}\delta^{\pm}(H)$ that holds for every $n$-vertex digraph $G$ with $\delta^{+}(G)\geq d$. In the regime when $d=\alpha n$ for some fixed $\alpha>0$, Theorem 1 guarantees the existence of a subgraph $H$ with $\frac{\delta^{\pm}(H)}{n}\geq\frac{\alpha^{2}}{2}$.
On the other hand, for any rational $\alpha<\frac{1}{3}$, let us fix $k:=\left\lfloor\frac{1}{\alpha}\right\rfloor-2$ and a large enough $d$ such that both $\ell:=\frac{d+1}{2k}$ and $n:=\frac{d}{\alpha}$ are integers. Note that for such a choice of the parameters, it holds that $(k+1)d+\ell<n$. These parameters satisfy the assumptions of Proposition 2, hence there exists an $n$-vertex tournament $T$ with $\delta^{+}(T)=d=\alpha n$ and $\max\limits_{H\subseteq G}\delta^{\pm}(H)\leq\ell$. In particular, if $\alpha\leq\frac{1}{4}$ then it holds that
$\frac{1}{n}\cdot{\max_{H\subseteq G}\delta^{\pm}(H)}\leq\frac{\alpha^{2}}{2}+6\alpha^{3}+O\left(\frac{1}{n}\right)\,,$ (2)
which is very close to the bound proven in Theorem 1 in case $\alpha$ is small.
4.1 Very dense digraphs
We have observed that if $\alpha>0$ is small, then the constant $\frac{\alpha^{2}}{2}$ in Theorem 1 is almost optimal; see (2). In this section, we prove that if $\alpha\geq 0.7832$ then every $n$-vertex digraph $G$ with $\delta^{+}(G)=\alpha n$ has $H\subseteq G$ with $\frac{\delta^{\pm}(H)}{n}>\frac{\alpha^{2}}{2}$.
Let us start with introducing some additional notation. Given an integer $n$ and $\alpha\in(0,1)$, let $\mathcal{G}_{n,\alpha}$ be the set of all $n$-vertex digraphs $G$ with $\delta^{+}(G)\geq\alpha n$. Next, let
$h(n,\alpha):=\min\limits_{G\in\mathcal{G}_{n,\alpha}}\max_{H\subseteq G}\frac{\delta^{\pm}(H)}{n}\hskip 20.00003pt\mbox{and}\hskip 20.00003pth(\alpha):=\liminf\limits_{n\to\infty}\,h(n,\alpha)\,.$
Note that $h(\alpha)\leq\alpha$ by considering random digraphs of density $\alpha$. Theorem 1 yields that $h(\alpha)\geq\frac{\alpha^{2}}{2}$ by repeatedly removing vertices of a given graph that have small indegree or outdegree. However, if $\alpha\geq 0.7832$ then one gets a stronger lower bound by removing only those vertices that have a small indegree. More precisely, we assert that
$h(\alpha)\geq 1-\sqrt{3-4\alpha+\alpha^{2}}\,,$
which is larger than $\frac{\alpha^{2}}{2}$ for $\alpha>\sqrt{2\sqrt{2}+2}-\sqrt{2}\sim 0.7832$.
To this end, fix an $n$-vertex digraph $G=(V,A)$. Without loss of generality, every vertex $v\in V$ satisfies $\deg_{G}^{+}(v)=\alpha n$. Now remove one by one all its vertices that currently have the indegree smaller than $\beta n$ for an appropriately chosen $\beta<\alpha$. This procedure terminates either when all the vertices of $G$ have been removed, or, when the minimum indegree of the current subgraph is at least $\beta n$.
Set $\beta:=1-\sqrt{3-4\alpha+\alpha^{2}}$. Suppose the indegree-greedy procedure described in the previous paragraph is applied to $G$, and let $\tau$ denotes the proportion of the vertices of $G$ that has been removed so far. Let $v_{1},v_{2},\ldots,v_{\tau n}$ be those vertices in the order of their removal, and, for every $i\in[\tau n]$ we set $W_{i}:=\{v_{1},v_{2},\ldots,v_{i}\}$. Counting the number of arcs of $G$ in two ways and rearranging yield that
$\small\alpha=\sum_{i=1}^{\tau n}\frac{|N_{G}^{-}(v_{i})\cap W_{i}|+|N_{G}^{-}(v_{i})\setminus W_{i}|}{n^{2}}+\sum_{v\in V\setminus W_{\tau n}}\frac{\deg_{G}^{-}(v)}{n^{2}}<\frac{\tau^{2}}{2}+\beta\tau+(1-\tau).$ (3)
Let $f(\tau):=\frac{\tau^{2}}{2}+\beta\tau+(1-\tau)-\alpha$. A bit tedious calculation reveals that $f(\alpha-\beta)=0$, which together with (3) implies that the procedure has to terminate after removing less than $(\alpha-\beta)n$ vertices from $G$.
Let $H\subseteq G$ be the remaining graph after the procedure terminates, and let $\tau_{0}$ be the total proportion of the removed vertices from $G$. In other words, $H$ has $(1-\tau_{0})n$ vertices. It follows that every vertex of $H$ has the indegree at least $\beta n$, and the outdegree at least $(\alpha-\tau_{0})n$. Therefore,
$\frac{\delta^{\pm}(H)}{n}\geq\min\{\alpha-\tau_{0},\beta\}\geq\min\{\alpha-(\alpha-\beta),\beta\}=\beta=1-\sqrt{3-4\alpha+\alpha^{2}}\,.$
In particular, it holds that $h(\alpha)\geq\max\left\{\frac{\alpha^{2}}{2},1-\sqrt{3-4\alpha+\alpha^{2}}\right\}$. We believe it is an interesting open problem to find stronger estimates on $h(\alpha)$.
Problem 1.
For a given $\alpha\in(0,1)$, determine the value of $h(\alpha)$.
4.2 Orientations of very dense undirected graphs
We say that a digraph $G$ is an oriented graph, i.e., an orientation of the edge-set of a simple undirected graph, if every pair of the vertices of $G$ is joined by at most one arc. For any $n$-vertex oriented graph $G$ with $\delta^{+}(G)=\alpha n$ it must hold that $\alpha<1/2$, so at the first glance the previous subsection does not provide anything useful for $G$. However, the double-counting argument (3) can be easily improved for oriented graphs since $\deg_{G}^{-}(v)<(1-\alpha)n$ for every vertex $v\in V(G)$. In the rest of this section, we show that this improvement yields the existence of a subgraph $H\subseteq G$ such that
$\frac{\delta^{\pm}(H)}{n}\geq 1-\alpha-\sqrt{3-8\alpha+4\alpha^{2}}\,.$ (4)
Observe that the right-hand side of (4) is larger than $\frac{\alpha^{2}}{2}$ for $\alpha>\alpha_{1}$, where $\alpha_{1}$ is the only positive real root of $x^{4}+4x^{3}-16x^{2}+24x-8$. We note that $\alpha_{1}<0.4528$.
Let us now introduce some additional notation, which will be analogous to the one in Section 4.1. For an integer $n$ and $\alpha\in(0,\frac{1}{2})$, let $\widehat{\mathcal{G}}_{n,\alpha}$ be the set of all $n$-vertex oriented graphs $G$ such that $\delta^{+}(G)\geq\alpha n$, and
[…]
Clearly, $\widehat{h}(\alpha)\geq h(\alpha)$, and $\widehat{h}(\alpha)\leq\alpha$ by considering, for example, random subsets of the arc-set of a balanced tournament.
We are now ready to explain the existence of $H\subseteq G$ satisfying (4). Fix an $n$-vertex oriented graph $G\in\widehat{\mathcal{G}}_{n,\alpha}$. Without loss of generality, we will again assume that $\deg_{G}^{+}(v)=\alpha n$ for every $v\in V(G)$. Now consider removing one by one all the vertices that have the current indegree smaller than $\widehat{\beta}n$ for some fixed $\widehat{\beta}>0$. As we have already mentioned, every vertex $v\in V(G)$ satisfies $\deg_{G}^{-}(v)<(1-\alpha)n$. Thus the double-counting argument (3) improves to the following:
$\alpha<\frac{\tau^{2}}{2}+\widehat{\beta}\tau+(1-\tau)(1-\alpha)\,.$ (5)
Let $H$ be the remaining subgraph after the indegree-greedy procedure terminates, and set
$\widehat{\beta}:=1-\alpha-\sqrt{3-8\alpha+4\alpha^{2}}\quad\mbox{and}\quad\widehat{f}(\tau):=\frac{\tau^{2}}{2}+\widehat{\beta}\tau+(1-\tau)(1-\alpha)-\alpha\,.$
We claim that $H$ satisfies (4). Indeed, since $\widehat{f}(\alpha-\widehat{\beta})=0$, the inequality (5) implies that $H$ has $(1-\tau_{0})n$ vertices for some $\tau_{0}<\alpha-\widehat{\beta}$. In particular,
$\frac{\delta^{\pm}(H)}{n}\geq\min\{\alpha-\tau_{0},\widehat{\beta}\}\geq\min\{\alpha-(\alpha-\widehat{\beta}),\widehat{\beta}\}=1-\alpha-\sqrt{3-8\alpha+4\alpha^{2}}\,.$
We have shown that $\widehat{h}(\alpha)\geq\max\left\{\frac{\alpha^{2}}{2},1-\alpha-\sqrt{3-8\alpha+4\alpha^{2}}\right\}$. As in the previous subsection, we find interesting to establish stronger estimates on $\widehat{h}(\alpha)$.
Problem 2.
For a given $\alpha\in\left(0,\frac{1}{2}\right)$, determine the value of $\widehat{h}(\alpha)$.
4.3 Normalized minimum semidegree
Throughout this subsection, fix a large $G\in\mathcal{G}_{n,\alpha}$. A variant of the question we consider in this paper is to maximize, over all subgraphs $H\subseteq G$, the value of $\delta^{\pm}(H)$ normalized by the number of vertices of $H$. In other words, our aim is to bound the expression
$\mu(G):=\max_{\emptyset\neq H\subseteq G}\frac{\delta^{\pm}(H)}{|V(H)|}\,.$ (6)
First observe that if $G$ is a balanced blow-up of a directed cycle of length $\left\lceil 1/\alpha\right\rceil$, then $\mu(G)\leq\alpha$. On the other hand, let us observe that $\mu(G)\geq\frac{\alpha}{2}$. The celebrated Caccetta-Häggkvist conjecture [3] states that $G$ contains a directed cycle $C$ of length at most $\left\lceil 1/\alpha\right\rceil$. In particular, $\mu(G)\geq\frac{1}{\left\lceil 1/\alpha\right\rceil}$. A result of Chvátal and Szemerédi [4] weakens the conclusion of the conjecture and establishes a directed cycle of length less than $\frac{2}{\alpha}$, thus a subgraph $H\subseteq G$ with $\frac{\delta^{\pm}(H)}{|V(H)|}\geq\frac{\alpha}{2}$ always exists. Note that the multiplicative constant $2$ from [4] was later improved by Shen [7].
The subgraph $H\subseteq G$ found in the previous paragraph is only of a constant order. It is natural to ask whether one can also find $H^{\prime}\subseteq G$ with $\frac{\delta^{\pm}(H^{\prime})}{|V(H^{\prime})|}=\Theta(\alpha)$ while $|V(H^{\prime})|$ is “large”. The answer is yes — a standard application of the regularity method for directed graphs (see, e.g., [1]) yields that $G\in\mathcal{G}_{n,\alpha}$ contains $\Theta\left(n^{\ell}\right)$ cycles of length $\ell$ for some $2\leq\ell\leq\frac{2}{\alpha-\varepsilon}$. Therefore, by supersaturation, $G$ also contains a balanced blow-up of $C_{\ell}$ with $\operatorname{polylog}(n)$ vertices.
Acknowledgements
We would like to thank Benny Sudakov for fruitful discussions on the topics related to this project.